### v1

@@ 4 Experiments
Figure 1: Robustness to varying levels of equal distortion: (Left) The estimated undistortion parameter and (Right) the pose error. An x-axis value of $-0.3$ means that we apply a distortion that can be undistorted with the undistortion parameter $-0.3$. The cyan line indicates the ground truth $\lambda$ value on the y-axis.
@@
Robustness to varying levels of distortion. In the first experiment, we study the robustness of different approaches w.r.t. increasing lens distortion. To this end, we distort the positions of the local features extracted from the undistorted images provided by the ETH3D dataset.444 In contrast to extracting feaures from images distorted with varying levels of distortion, this approach isolates the effect of distortion on the different strategies since the same feature matches are used for each distortion level. We apply the same amount of distortion to both images in each image pair using the inverse of  (2) (see the SM for the closed-form formula for the inverse). The added distortion ranges from 0 to the distortion that corresponds to the undistortion parameter $\lambda=-1.8$. This represents scenarios from no/small to significant distortion. For each distortion, we run GC-RANSAC on all image pairs with different combinations of minimal and non-minimal solvers, including the three versions of the 7pt solver with sampled parameters: (1) $\mathbf{U}=\{0\}$, (2) $\mathbf{U}=\{-0.9\}$, and (3) $\mathbf{U}=\{0,-0.6,-1.2\}$. Fig. 1 shows the estimated undistortion parameter $\lambda$ (left) and the pose error (right). The rotation and translation error, as well as results for one scene from the Phototourism dataset are in the SM. Naturally, estimating the distortion already in the first step of RANSAC (8pt $\mathbf{F}\lambda$+ 9pt $\mathbf{F}\lambda$ and 9pt $\mathbf{F}\lambda$+ 9pt $\mathbf{F}\lambda$) leads to the best results. Still, sampling-based strategies followed by LO with 9pt $\mathbf{F}\lambda$ can perform similarly well if they use samples close enough to the true level of distortion. For three samples, slightly larger errors were observed only for larger distortion values ($\lambda\leq-1.5$). However, this can be easily eliminated by adding one more sample and running the 7pt solver four times.
@@
Minimal Refinement Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 29.45 15.46 0.14 0.24 0.37 0.87 0.86 111.85
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ 00 11.43 13.32 0.44 0.57 0.69 0.14 0.05 143.07
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 00 14.48 14.04 0.40 0.53 0.65 0.57 0.11 145.54
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 33.51 20.14 0.09 0.17 0.29 0.87 0.87 112.61
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 00 15.16 14.17 0.39 0.52 0.64 0.60 0.11 147.99
@@
Minimal Refinement Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 12.91 4.47 0.37 0.51 0.65 0.15 0.15 111.22
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ 00 10.74 3.15 0.45 0.58 0.71 0.12 0.04 143.62
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 00 13.45 3.67 0.42 0.55 0.68 0.50 0.09 146.01
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 13.95 5.14 0.33 0.47 0.62 0.15 0.15 112.56
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 00 13.58 3.71 0.41 0.55 0.67 0.51 0.10 150.26
@@
Minimal Refinement Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 34.85 21.10 0.08 0.16 0.28 1.10 1.10 112.23
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ $-0.9$ 10.75 13.18 0.45 0.58 0.70 0.13 0.05 142.14
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ $-0.9$ 13.43 13.78 0.41 0.54 0.67 0.50 0.11 144.73
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 36.05 22.65 0.07 0.14 0.26 1.10 1.10 112.69
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ $-0.9$ 14.14 13.94 0.40 0.53 0.66 0.56 0.11 147.27
@@
Scenario A - Wild: In the first scenario, we assume no knowledge about the cameras. The cameras can have distortions ranging from small to very high. This scenario represents, e.g., images downloaded from the Internet. To simulate this scenario, we distorted the positions of the local features extracted from the undistorted images provided by the ETH3D and Phototourism datasets. For each pair of images, we sample undistortion parameters from a distribution $\mathcal{U}$. Given an undistortion parameter $\lambda$, we apply the corresponding distortion to an image using the inverse function of (2). We apply either the same or different distortions based on the studied setup ($\mathbf{F}\lambda$ or $\mathbf{F}\lambda_{1}\lambda_{2}$). We define $\mathcal{U}$ as a piecewise distribution, which is uniform between $-1.5$ and 00, while its density decreases linearly from $-1.5$ to $-1.8$, reaching half the density of the uniform range. This is done to simulate that in practice, distortion parameters in the range $[-1.5,0]$ are more common than in the range $[-1.8,-1.5]$. Thus, it is natural to sample undistortion parameters from a wide rage of parameters for the sampling-based approach. Based on the results of the solver with $\mathbf{U}_{1}=\mathbf{U}_{2}=\{0,-0.6,-1.2\}$ in Fig. 1, we evaluated this solver for this scenario. Tab. 1 (top) shows the results for the Phototourism datasets and Poselib RANSAC. Results for this scenario for ETH3D and for GC-RANSAC are shown in the SM. Poselib RANSAC does not use non-minimal solvers in LO but refines poses using LM. Thus, the Refinement column indicates which parameters are optimized inside LO. We show results for cameras with the same distortion ($\lambda_{1}=\lambda_{2})$ and for cameras with different distortions ($\lambda_{1}\not=\lambda_{2})$. In this scenario, we also tested variants in which we optimize different distortions even for cameras with the same distortion. As can be seen in Tab. 1, the 7pt solver with the sampling-based strategy with $\mathbf{U}_{1}=\mathbf{U}_{2}=\{0,-0.6,-1.2\}$ even outperforms the dedicated minimal radial distortion solvers that are applied in the first step of RANSAC. The non-minimal 9pt $\mathbf{F}\lambda$ and 10pt $\mathbf{F}\lambda_{1}\lambda_{2}$ solvers are faster than the sampling-based strategy. Yet, they are significantly less accurate.
@@
Scenario B - Small distortion: In the second scenario, we simulate prior knowledge that our cameras have small distortion, e.g., we are processing images taken by mobile phone or DSLR cameras. To simulate this scenario, we distort the feature points with distortions corresponding to undistortion parameters uniformly sampled from the interval $[-0.3,0]$. In this case, it makes sense to run the 7pt solver in the sampling-based strategy only once with a small undistortion parameter. We decided to use $\mathbf{U}_{1}=\mathbf{U}_{2}=\{0\}$ to simulate the standard baseline. Tab. 1 (middle) shows the results for this scenario. In this case, it can be seen that the baseline 7pt solver followed by LO of the radial distortion parameters outperforms the minimal radial distortion solvers at competitive run-times.
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 17.17 16.62 0.21 0.36 0.52 0.53 0.09 1420.61
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 28.70 18.33 0.24 0.35 0.48 0.55 0.09 1350.41
@@
Minimal Refinement Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 21.34 6.58 0.29 0.42 0.56 0.53 0.09 1125.96
@@
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ 00 18.45 4.36 0.38 0.51 0.63 0.35 0.10 1124.65
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 00 18.66 4.22 0.39 0.52 0.64 0.37 0.11 1124.28
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 35.31 8.00 0.31 0.41 0.51 0.55 0.09 1125.09
@@
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 00 32.90 4.76 0.39 0.49 0.57 0.44 0.18 1107.31
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 11.16 5.48 0.25 0.40 0.58 0.69 0.72 1620.91
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 16.71 9.51 0.14 0.27 0.45 0.64 0.68 1240.02
@@
Minimal Refinement Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 13.12 4.83 0.38 0.50 0.65 0.69 0.72 1128.75
@@
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ 00 17.40 1.11 0.68 0.77 0.84 0.19 0.05 1112.58
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 00 17.38 1.19 0.67 0.76 0.84 0.21 0.06 1110.99
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 18.13 8.72 0.23 0.35 0.52 0.64 0.68 1145.52
@@
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 00 19.47 1.84 0.57 0.67 0.77 0.25 0.11 1177.30
@@
Scenario C - Visible distortion: In the last scenario, we assume that we know that our images have visible (but unknown) distortion. To simulate this, we distort the feature points with distortions corresponding to undistortion parameters uniformly sampled from the interval $[-1.8,-0.5]$. For the sampling-based strategy, we tested two different variants: (1) $\mathbf{U}_{1}=\mathbf{U}_{2}=\{-0.9\}$ and (2) $\mathbf{U}_{1}=\mathbf{U}_{2}=\{-0.6,-0.9,-1.2\}$. The results are shown in Tab. 1 (bottom). Again, the 7pt solver with $\mathbf{U}_{1}=\mathbf{U}_{2}=\{-0.6,-0.9,-1.2\}$ outperforms the minimal radial distortion solvers. Even sampling one distortion -0.9 leads to similar results to the 8pt $\mathbf{F}\lambda$ solver, while being much faster. For the case $\lambda_{1}\not=\lambda_{2}$, both sampling strategies significantly outperform the radial distortion solvers.

### v2

@@ 4 Experiments
Figure 1: Robustness to varying levels of equal distortion: (Left) The estimated undistortion parameter and (Right) the pose error. An x-axis value of 0.3-0.3 means that we apply a distortion that can be undistorted with the undistortion parameter 0.3-0.3. The cyan line indicates the ground truth $\lambda$ value on the y-axis.
@@
Robustness to varying levels of distortion. In the first experiment, we study the robustness of different approaches w.r.t. increasing lens distortion. To this end, we distort the positions of the local features extracted from the undistorted images provided by the ETH3D dataset. [^4] In contrast to extracting feaures from images distorted with varying levels of distortion, this approach isolates the effect of distortion on the different strategies since the same feature matches are used for each distortion level. We apply the same amount of distortion to both images in each image pair using the inverse of  (2) (see the SM for the closed-form formula for the inverse). The added distortion ranges from 0 to the distortion that corresponds to the undistortion parameter $\lambda=-1.8$. This represents scenarios from no/small to significant distortion. For each distortion, we run GC-RANSAC on all image pairs with different combinations of minimal and non-minimal solvers, including the three versions of the 7pt solver with sampled parameters: (1) $\mathbf{U}=\{0\}$, (2) $\mathbf{U}=\{-0.9\}$, and (3) $\mathbf{U}=\{0,-0.6,-1.2\}$. Fig. 1 shows the estimated undistortion parameter $\lambda$ (left) and the pose error (right). The rotation and translation error, as well as results for one scene from the Phototourism dataset are in the SM. Naturally, estimating the distortion already in the first step of RANSAC (8pt $\mathbf{F}\lambda$+ 9pt $\mathbf{F}\lambda$ and 9pt $\mathbf{F}\lambda$+ 9pt $\mathbf{F}\lambda$) leads to the best results. Still, sampling-based strategies followed by LO with 9pt $\mathbf{F}\lambda$ can perform similarly well if they use samples close enough to the true level of distortion. For three samples, slightly larger errors were observed only for larger distortion values ($\lambda\leq-1.5$). However, this can be easily eliminated by adding one more sample and running the 7pt solver four times.
@@
Minimal Refinement Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 29.45 15.46 0.14 0.24 0.37 0.87 0.86 111.85
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ 0 11.43 13.32 0.44 0.57 0.69 0.14 0.05 143.07
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0 14.48 14.04 0.40 0.53 0.65 0.57 0.11 145.54
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 33.51 20.14 0.09 0.17 0.29 0.87 0.87 112.61
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0 15.16 14.17 0.39 0.52 0.64 0.60 0.11 147.99
@@
Minimal Refinement Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 12.91 4.47 0.37 0.51 0.65 0.15 0.15 111.22
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ 0 10.74 3.15 0.45 0.58 0.71 0.12 0.04 143.62
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0 13.45 3.67 0.42 0.55 0.68 0.50 0.09 146.01
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 13.95 5.14 0.33 0.47 0.62 0.15 0.15 112.56
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0 13.58 3.71 0.41 0.55 0.67 0.51 0.10 150.26
@@
Minimal Refinement Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 34.85 21.10 0.08 0.16 0.28 1.10 1.10 112.23
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ 0.9-0.9 10.75 13.18 0.45 0.58 0.70 0.13 0.05 142.14
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0.9-0.9 13.43 13.78 0.41 0.54 0.67 0.50 0.11 144.73
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 36.05 22.65 0.07 0.14 0.26 1.10 1.10 112.69
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0.9-0.9 14.14 13.94 0.40 0.53 0.66 0.56 0.11 147.27
@@
Scenario A - Wild: In the first scenario, we assume no knowledge about the cameras. The cameras can have distortions ranging from small to very high. This scenario represents, e.g., images downloaded from the Internet. To simulate this scenario, we distorted the positions of the local features extracted from the undistorted images provided by the ETH3D and Phototourism datasets. For each pair of images, we sample undistortion parameters from a distribution $\mathcal{U}$. Given an undistortion parameter $\lambda$, we apply the corresponding distortion to an image using the inverse function of (2). We apply either the same or different distortions based on the studied setup ($\mathbf{F}\lambda$ or $\mathbf{F}\lambda_{1}\lambda_{2}$). We define $\mathcal{U}$ as a piecewise distribution, which is uniform between 1.5-1.5 and 0, while its density decreases linearly from 1.5-1.5 to 1.8-1.8, reaching half the density of the uniform range. This is done to simulate that in practice, distortion parameters in the range [1.5,0][-1.5,0] are more common than in the range [1.8,1.5][-1.8,-1.5]. Thus, it is natural to sample undistortion parameters from a wide rage of parameters for the sampling-based approach. Based on the results of the solver with $\mathbf{U}_{1}=\mathbf{U}_{2}=\{0,-0.6,-1.2\}$ in Fig. 1, we evaluated this solver for this scenario. Tab. 1 (top) shows the results for the Phototourism datasets and Poselib RANSAC. Results for this scenario for ETH3D and for GC-RANSAC are shown in the SM. Poselib RANSAC does not use non-minimal solvers in LO but refines poses using LM. Thus, the Refinement column indicates which parameters are optimized inside LO. We show results for cameras with the same distortion ($\lambda_{1}=\lambda_{2})$ and for cameras with different distortions ($\lambda_{1}\not=\lambda_{2})$. In this scenario, we also tested variants in which we optimize different distortions even for cameras with the same distortion. As can be seen in Tab. 1, the 7pt solver with the sampling-based strategy with $\mathbf{U}_{1}=\mathbf{U}_{2}=\{0,-0.6,-1.2\}$ even outperforms the dedicated minimal radial distortion solvers that are applied in the first step of RANSAC. The non-minimal 9pt $\mathbf{F}\lambda$ and 10pt $\mathbf{F}\lambda_{1}\lambda_{2}$ solvers are faster than the sampling-based strategy. Yet, they are significantly less accurate.
@@
Scenario B - Small distortion: In the second scenario, we simulate prior knowledge that our cameras have small distortion, e.g., we are processing images taken by mobile phone or DSLR cameras. To simulate this scenario, we distort the feature points with distortions corresponding to undistortion parameters uniformly sampled from the interval [0.3,0][-0.3,0]. In this case, it makes sense to run the 7pt solver in the sampling-based strategy only once with a small undistortion parameter. We decided to use $\mathbf{U}_{1}=\mathbf{U}_{2}=\{0\}$ to simulate the standard baseline. Tab. 1 (middle) shows the results for this scenario. In this case, it can be seen that the baseline 7pt solver followed by LO of the radial distortion parameters outperforms the minimal radial distortion solvers at competitive run-times.
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 17.17 16.62 0.21 0.36 0.52 0.53 0.09 1420.61
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 28.70 18.33 0.24 0.35 0.48 0.55 0.09 1350.41
@@
Minimal Refinement Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 21.34 6.58 0.29 0.42 0.56 0.53 0.09 1125.96
@@
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ 0 18.45 4.36 0.38 0.51 0.63 0.35 0.10 1124.65
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0 18.66 4.22 0.39 0.52 0.64 0.37 0.11 1124.28
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 35.31 8.00 0.31 0.41 0.51 0.55 0.09 1125.09
@@
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0 32.90 4.76 0.39 0.49 0.57 0.44 0.18 1107.31
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 11.16 5.48 0.25 0.40 0.58 0.69 0.72 1620.91
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 16.71 9.51 0.14 0.27 0.45 0.64 0.68 1240.02
@@
Minimal Refinement Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 13.12 4.83 0.38 0.50 0.65 0.69 0.72 1128.75
@@
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ 0 17.40 1.11 0.68 0.77 0.84 0.19 0.05 1112.58
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0 17.38 1.19 0.67 0.76 0.84 0.21 0.06 1110.99
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 18.13 8.72 0.23 0.35 0.52 0.64 0.68 1145.52
@@
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0 19.47 1.84 0.57 0.67 0.77 0.25 0.11 1177.30
@@
Scenario C - Visible distortion: In the last scenario, we assume that we know that our images have visible (but unknown) distortion. To simulate this, we distort the feature points with distortions corresponding to undistortion parameters uniformly sampled from the interval [1.8,0.5][-1.8,-0.5]. For the sampling-based strategy, we tested two different variants: (1) $\mathbf{U}_{1}=\mathbf{U}_{2}=\{-0.9\}$ and (2) $\mathbf{U}_{1}=\mathbf{U}_{2}=\{-0.6,-0.9,-1.2\}$. The results are shown in Tab. 1 (bottom). Again, the 7pt solver with $\mathbf{U}_{1}=\mathbf{U}_{2}=\{-0.6,-0.9,-1.2\}$ outperforms the minimal radial distortion solvers. Even sampling one distortion -0.9 leads to similar results to the 8pt $\mathbf{F}\lambda$ solver, while being much faster. For the case $\lambda_{1}\not=\lambda_{2}$, both sampling strategies significantly outperform the radial distortion solvers.